# Topic-FlipRAG: Topic-Orientated Adversarial Opinion Manipulation Attacks to Retrieval-Augmented Generation Models

### v2

@@ Abstract.
Retrieval-Augmented Generation (RAG) systems based on Large Language Models (LLMs) have become essential for tasks such as question answering and content generation. However, their increasing impact on public opinion and information dissemination has made them a critical focus for security research due to inherent vulnerabilities. Previous studies have predominantly addressed attacks targeting factual or single-query manipulations. In this paper, we address a more practical scenario: topic-oriented adversarial opinion manipulation attacks on RAG models, where LLMs are required to reason and synthesize multiple perspectives, rendering them particularly susceptible to systematic knowledge poisoning. Specifically, we propose Topic-FlipRAG, a two-stage manipulation attack pipeline that strategically crafts adversarial perturbations to influence opinions across related queries. This approach combines traditional adversarial ranking attack techniques and leverages the extensive internal relevant knowledge and reasoning capabilities of LLMs to execute semantic-level perturbations. Experiments show that the proposed attacks effectively shift the opinion of the model’s outputs on specific topics, significantly impacting user information perception. Current mitigation methods cannot effectively defend against such attacks, highlighting the necessity for enhanced safeguards for RAG systems, and offering crucial insights for LLM security research.
Corresponding author: laujames2017@whu.edu.cn
1. Introduction
@@
Figure 1. An concise overview of topic-oriented adversarial opinion manipulation attacks on RAG systems.
@@
Retrieval-Augmented Generation (RAG) systems, built on large language models (LLMs), have advanced significantly and been widely applied in tasks like question answering and content generation (Gao et al., 2023; Zhao et al., 2024). These systems integrate retrieval mechanisms with generative models, accessing diverse sources, e.g., Wikipedia, Reddit, and news articles, to deliver updated knowledge. While research has focused on the development of RAG frameworks (Asai et al., 2023; Shi et al., 2025; Xu et al., 2024), the majority of studies emphasize performance and generalization. Consequently, security aspects have received comparatively little attention.
@@
A defining feature of RAG systems is the vast scale of their referenced corpora or knowledge bases. However, these documents might originate from sources beyond service providers’ control and may resist complete content purification. This creates new security risks for RAG-supported LLMs: adversaries can inject meticulously crafted malicious content (Carlini et al., 2024) into retrieval collections, ensuring its prioritized retrieval to influence LLM outputs. The inherent vulnerabilities of RAG models, particularly their persistent impact on public opinion and information dissemination (Chen et al., 2025), have gradually drawn significant attention from the AI safety research community (Zhong et al., 2023; Xue et al., 2024; Zou et al., 2024). Investigating these vulnerabilities is critical for advancing RAG security and ensuring system robustness and credibility.
@@
Early adversarial attacks on RAG systems primarily focused on jailbreak techniques and query-specific perturbations (Chen et al., 2025; Zhong et al., 2023). These methods tamper with knowledge bases to manipulate retrieval rankings, enabling LLMs to provide answers defined by attackers for specific queries. Recent research, such as PoisonedRAG (Zou et al., 2024), illustrates how poisoned texts can lead to the generation of misinformation, revealing critical vulnerabilities in applications like cybersecurity and healthcare. However, these attacks often lack practicality as they concentrate on isolated factual queries. Such queries can often be countered through mitigation strategies such as reranking (De Stefano et al., 2024), which tend to overlook thematic or topic-level manipulation. Additionally, existing studies largely overlook controversial topics, where limited user understanding may increase susceptibility to opinion manipulation.
This paper addresses this critical gap by concentrating on the nuanced manipulation of opinions through adversarial attacks targeted at topics within RAG models, which presents a novel and urgent challenge given the increasing complexity and reliance on these systems.
@@
In line with settings from previous data poisoning findings(Zou et al., 2024; Carlini et al., 2024), we consider a scenario where the attacker could inject a few carefully crafted poisoned texts into the knowledge base. For example, if the knowledge base includes texts sourced from Wikipedia, the adversary might inject poisoned content by maliciously editing Wikipedia entries. Besides, we focus on more practical and challenging black-box scenario, where no model information is disclosed, except that the attackers can query the target RAG and obtain responses, which contains the corresponding candidate referential documents (Chen et al., 2025).
@@
In this paper, we propose the Topic-FlipRAG, which is a two-stage multi-granularity attack method designed to manipulating the stance polarity of RAGs with camouflaged modification of target documents. Specifically, in the first stage, we perform a stealthy adversarial modifications on target documents. This is achieved by incorporating the extensive internal general semantic knowledge in LLMs with its analysis and reasoning ability. In the second stage, inspired by adversarial semantic collision (Song et al., 2020), we utilize gradients from an open-sourced neural ranking model (NRM) to generate topic-specific adversarial triggers.
@@
Our experiments show Topic-FlipRAG achieves 0.5 average stance variation (ASV) across four domains, significantly outperforming other baselines. It demonstrates such adversarial attacks can effectively manipulate RAG models to produce outputs aligned with specific topical stances, thus influencing how information is presented and perceived. Based on that, we also conduct the user experiments, which reveal a significant impact on users’ polarities toward controversial topics, with polarity shifts exceeding 16% after interacting with poisoned RAG systems. Futhermore, we explore several potential mitigation strategy, including perplexity-based detection, random masking, and reranking. Our results show these mitigations are inadequate against adversarial opinion manipulation attack, especially for Topic-FlipRAG, underscoring the necessity for novel mitigation strategies.
@@
(3) Through extensive experiments, we demonstrate that Topic-FlipRAG achieves significantly better attack performance across all metrics compared to baseline methods. Furthermore, we evaluate its practical impact through user experiments, highlighting its potential to manipulate user opinions.
@@
(4) We analyze the efficacy of several defense mechanisms and demonstrate their insufficiency in countering Topic-FlipRAG. This underscores the urgent need for robust and adaptive defense strategies against such attacks.
2. Background and Related Work
2.1. Retrieval Augmented Generation (RAG)
@@
RAG models enhance their responses by accessing and incorporating external knowledge from large-scale databases or corpora during the generation process (Lewis et al., 2020; Shi et al., 2025; Liu et al., 2023a). By leveraging external data sources, RAG can provide more accurate and comprehensive answers, especially for queries requiring up-to-date information or specialized knowledge that may not be well-represented in the model’s training data (Siyue et al., 2024). Moreover, it can scale more effectively and flexibly by updating the retrieval corpus without necessitating extensive retraining of the generative component (Gao et al., 2023; Wu et al., 2024).
@@
The workflow of a RAG systems is generally divided into two sequential phases: retrieval and generation. In the retrieval phase, upon the submission of a user query $q$, the RAG system retrieves $k$ relevant documents from the corpus $D$ with the highest embedding similarities to the query $q$. Specifically, for each document $d\in D$, the relevance score with the query $q$ is computed as $R(q,d)$. In this paper, we adopt the a more practical and widely used framework in LangChain that the retrieval phase is incorporated with historical-aware query rewriting and intention reasoning. In the generation phase, Given a the rewritten query $q^{\prime}$ , a set of top-k retrieved documents $D_{k}$, and access to the LLM, one can query the LLM with $q^{\prime}$ with the top-k retrieved documents $D_{k}$, the LLM generates an answer for original query $q$ by leveraging $D_{k}$ as context. This process is detailed in Appendix A.1.
2.2. Attacks to Retriever
@@
Retriever vulnerabilities in RAG systems pose significant security risks due to their neural ranking dependencies (Goren et al., 2018). Adversarial attacks manipulate document rankings through semantic perturbations to promote target documents in query-specific results (Liu et al., 2022; Wu et al., 2023), undermining RAG’s core assumption of reliable high-ranked contexts. Current attacks are characterized along two axes: (1) Knowledge accessibility: White-box (full model access) vs. black-box (query-only) approaches. (2) Perturbation granularity: Word-level (Raval and Verma, 2020; Liu et al., 2023b, c; Wu et al., 2023), Phrase-level (Song et al., 2020; Liu et al., 2022; Chen et al., 2025), Sentence-level (Chen et al., 2023), and Hybrid (Liu et al., 2024).These neural semantic attacks parallel black-hat SEO tactics (Gyongyi and Garcia-Molina, 2005) but introduce novel challenges. Successful attacks propagate adversarial content to generators, enabling misinformation injection while evading traditional safeguards (Zhong et al., 2023; Chen et al., 2025), necessitating integrated security frameworks for RAG systems.
2.3. Attacks to LLMs
@@
Existing attacks on LLMs include jailbreak attacks (Wei et al., 2024; Deng et al., 2024; Li et al., 2023; Lin et al., 2024), backdoor attacks (Kandpal et al., 2023; Lu et al., 2024; Cao et al., 2023), prompt injection (Liu et al., 2023; Perez and Ribeiro, 2022; Greshake et al., 2023), and poisoning attacks (Shafran et al., 2024; Zou et al., 2024; Cho et al., 2024; Zhong et al., 2023; Chen et al., 2025; Zhang et al., 2024). Poisoning attacks uniquely threaten RAG systems by injecting adversarial content into retrieval corpora to manipulate outputs. Current RAG poisoning studies focus on closed-domain factoid QA pairs (e.g., ”CEO of OpenAI”) (Zou et al., 2024; Zhong et al., 2023; Cho et al., 2024; Shafran et al., 2024; Zhang et al., 2024), overlooking the multi-query nature of real-world topic exploration(Liu et al., 2023c; Xue et al., 2024), e.g., ”smartwatch battery life” and ”health tracking accuracy” under the ”wearable tech” theme. This motivates our focus on practical topic-level universal perturbations mirroring universal adversarial examples.
@@
Furthermore, existing defenses based on fact-checking are insufficient for opinion-based queries, such as ”Should genetic testing be regulated?”, which require nuanced reasoning rather than simple factual recall. Building on the insights from FlipedRAG (Chen et al., 2025), we tackle this significant gap by examining the adversarial manipulation of controversial topics. In such cases, LLMs are required to reason and synthesize multiple perspectives, making them particularly vulnerable to systematic knowledge poisoning.
3. Threat Model
@@
Given a set of topic-queries $\mathbf{Q}=\{q_{1},q_{2},\ldots,q_{|\mathbf{Q}|}\}$, a RAG corpus $D$, and a desired target polarity $S_{t}$ (e.g., Pro or Con), The adversary’s objective is to steer the RAG system’s overall stance across the entire set of queries toward stance $S_{t}$. Let $\text{doc}_{\text{tar}}$ be the target document the attacker aims to promote. The black-box retrieval model RM assigns a relevance score $R(q,d)$ to each document $d\in D$ and returns the top-$k$ documents as $\text{RM}_{k}(q)=\{d_{1},d_{2},\ldots,d_{k}\}$. The LLM then generates an answer $(q,\text{RM}_{k}(q))$, and the extracted opinion from this response is $S_{o}=\text{S}(\text{LLM}(q,RM_{k}(q)))$.
@@
Figure 2. An overview of our proposed Topic-Orientated Adversarial Opinion Manipulation Attack method for RAG systems.
3.1. Objective of the Adversary
@@
The adversary seeks to modify the target document $doc_{tar}$ through a series of subtle perturbations $p_{adv}$, while preserving its original semantics and fluency, thereby transforming the corpus $D$ into $D^{\prime}=D(\text{doc};\text{doc}_{\text{tar}}\oplus p_{\text{adv}})$. By increasing the retrieval model’s relevance score $R(q,\text{doc}_{\text{tar}}\oplus p_{\text{adv}})$ for each query $q\in\mathbf{Q}$, the modified document becomes more likely to appear among the top-$k$ results $\text{RM}_{k}(q)$.
@@
As a result, when the LLM generates answers conditioned on $(q,\text{RM}_{k}(q))$, the adversarially augmented document will influence the opinion $S(\text{LLM}(q,\text{RM}_{k}(q)))$ towards the target polarity $S_{t}$. Crucially, the objective extends beyond a single query: the adversary aims to ensure that, on average, the opinions produced by the LLM across the entire query set $\mathbf{Q}$ align with $S_{t}$. In other words, the attacker seeks:
$\max_{p_{\text{adv}}}\frac{1}{|\mathbf{Q}|}\sum_{q\in\mathbf{Q}}I\bigl{(}S(\text{LLM}(q,\text{RM}_{k}(q)))=S_{t}\bigr{)}\hskip 14.22636pt$
@@
where $I(\cdot)$ is an indicator function that returns $1$ if the generated stance matches $S_{t}$ and 00 otherwise. By optimizing for this criterion, the adversary ensures that the aggregated stance across all topic queries shifts significantly toward the desired polarity.
3.2. Capabilities of the Adversary

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) systems based on Large Language Models (LLMs) have become essential for tasks such as question answering and content generation. However, their increasing impact on public opinion and information dissemination has made them a critical focus for security research due to inherent vulnerabilities. Previous studies have predominantly addressed attacks targeting factual or single-query manipulations. In this paper, we address a more practical scenario: topic-oriented adversarial opinion manipulation attacks on RAG models, where LLMs are required to reason and synthesize multiple perspectives, rendering them particularly susceptible to systematic knowledge poisoning. Specifically, we propose Topic-FlipRAG, a two-stage manipulation attack pipeline that strategically crafts adversarial perturbations to influence opinions across related queries. This approach combines traditional adversarial ranking attack techniques and leverages the extensive internal relevant knowledge and reasoning capabilities of LLMs to execute semantic-level perturbations. Experiments show that the proposed attacks effectively shift the opinion of the model’s outputs on specific topics, significantly impacting users’ information perception. Current mitigation methods cannot effectively defend against such attacks, highlighting the necessity for enhanced safeguards for RAG systems, and offering crucial insights for LLM security research.
1 Introduction
@@
Figure 1: A concise overview of topic-oriented adversarial opinion manipulation attacks on RAG systems.
@@
Retrieval-Augmented Generation (RAG) systems, built on large language models (LLMs), have advanced significantly and been widely applied in tasks like question answering and content generation [gao2023retrieval, zhao2024retrieval]. These systems integrate retrieval mechanisms with generative models, accessing diverse sources, e.g., Wikipedia, Reddit, and news articles, to deliver updated knowledge. While research has focused on RAG frameworks [asai2023self, shi2025know, xu2024search], the majority of studies emphasize performance and generalization. Consequently, security aspects have received comparatively little attention.
@@
A defining feature of RAG systems is the vast scale of their referenced corpora or knowledge bases. However, these documents might originate from sources beyond service providers’ control and may resist complete content purification. This creates new security risks for RAG-supported LLMs: adversaries can inject meticulously crafted malicious content [carlini2024poisoning] into retrieval collections, ensuring its prioritized retrieval to influence LLM outputs. The inherent vulnerabilities of RAG models common to neural architectures, particularly their persistent impact on public opinion and information dissemination[chen2025flipedrag, zhong2023poisoning, chen2024research], have gradually drawn significant attention from the AI safety research community [zhong2023poisoning, xue2024badrag, zou2024poisonedrag]. Investigating these vulnerabilities is critical for advancing RAG security and ensuring system robustness and credibility.
@@
Early adversarial attacks on RAG systems primarily focused on jailbreak techniques and query-specific perturbations [chen2025flipedrag, zhong2023poisoning]. These methods tamper with knowledge bases to manipulate retrieval rankings, enabling LLMs to provide answers defined by attackers for specific query-level, such as PoisonedRAG [zou2024poisonedrag], revealing critical vulnerabilities in applications like cybersecurity and healthcare. However, these attacks often lack practicality as they concentrate on isolated factual queries. Such attacks can often be countered through mitigation strategies such as reranking and filtering [de2024rag, liu2023topic]. In contrast, thematic or topic-level attacks involve manipulating the overall perspective or stance of model outputs on broader queries, as shown in Figure 1. This form of manipulation is more practical and difficult to mitigate. Additionally, previous works overlook controversial topics, where limited user understanding may increase susceptibility to opinion manipulation.
@@
This paper addresses this critical gap by concentrating on the nuanced manipulation of opinions through adversarial attacks targeted at topics within RAG models, which presents a novel and urgent challenge given the increasing complexity and reliance on these systems. In line with settings from previous data poisoning findings[zou2024poisonedrag, carlini2024poisoning], we consider a scenario where the attacker could inject a few carefully crafted poisoned texts into the knowledge base. For example, if the knowledge base includes texts sourced from Wikipedia, the adversary might inject poisoned content by maliciously editing Wikipedia entries. Besides, we focus on a more practical and challenging black-box scenario, where no model information is disclosed, except that the attackers can query the target RAG and obtain responses, which contains the corresponding candidate referential documents [chen2025flipedrag].
@@
In this paper, we propose Topic-FlipRAG, which is a two-stage multi-granularity attack method designed to manipulating the stance polarity of RAGs with camouflaged modification of target documents. Specifically, in the first stage, we perform a stealthy adversarial modifications on target documents. This is achieved by incorporating the extensive internal general semantic knowledge in LLMs with its analysis and reasoning ability. In the second stage, inspired by adversarial semantic collision [song2020adversarial], we utilize gradients from an open-sourced neural ranking model (NRM) to generate topic-specific adversarial triggers.
@@
Our experiments show Topic-FlipRAG achieves 0.5 average stance variation (ASV) across four domains, significantly outperforming other baselines. It demonstrates such adversarial attacks can effectively manipulate RAG models to produce outputs aligned with specific topical stances, thus influencing how information is presented and perceived. Based on that, we also conduct the user experiments, which reveal a significant impact on users’ opinion polarities toward controversial topics, with polarity shifts exceeding 16% after interacting with poisoned RAG systems. Futhermore, we explore several potential mitigation strategies, including perplexity-based detection, random masking, paraphrasing and reranking. Our results show these mitigations are inadequate against adversarial opinion manipulation attack, especially for Topic-FlipRAG, underscoring the necessity for novel mitigation strategies.
@@
(3) Through extensive experiments, we demonstrate that Topic-FlipRAG significantly outperforms baseline methods across all evaluation metrics. Moreover, user studies reveal its substantial practical impact, highlighting its potential to manipulate user opinions.
@@
(4) We analyze the effectiveness of several existing defense mechanisms and demonstrate their inadequacy in mitigating attacks by Topic-FlipRAG. These findings highlight the urgent need for more robust and adaptive defense strategies.
2 Background and Related Work
2.1 Retrieval Augmented Generation (RAG)
@@
RAG models enhance their responses by accessing and incorporating external knowledge from large-scale databases or corpora during the generation process [lewis2020retrieval, shi2025know, liu2023webglm]. By leveraging external data sources, RAG can provide more accurate and comprehensive answers, especially for queries requiring up-to-date information or specialized knowledge that may not be well-represented in the model’s training data [siyue2024mrag]. Moreover, it can scale more effectively and flexibly by updating the retrieval corpus without necessitating extensive retraining of the generative component [gao2023retrieval, wu2024retrieval].
@@
The workflow of a RAG systems is generally divided into two sequential phases: retrieval and generation. In the retrieval phase, upon the submission of a user query $q$, the RAG system retrieves $k$ relevant documents from the corpus $D$ with the highest embedding similarities to the query $q$. Specifically, for each document $d\in D$, the relevance score with the query $q$ is computed as $R(q,d)$. In this paper, we adopt a more practical and widely used framework in LangChain that the retrieval phase is incorporated with historical-aware query rewriting and intention reasoning. In the generation phase, Given a rewritten query $q^{\prime}$ , a set of top-k retrieved documents $D_{k}$, and access to the LLM, one can query the LLM with $q^{\prime}$ with the top-k retrieved documents $D_{k}$, the LLM generates an answer for original query $q$ by leveraging $D_{k}$ as context.
2.2 Attacks to Retriever
@@
Retriever vulnerabilities in RAG systems pose significant security risks due to their neural ranking dependencies [goren2018ranking]. Adversarial attacks manipulate document rankings through semantic perturbations to promote target documents in query-specific results [liu2022order, wu2023prada], undermining RAG’s core assumption of reliable high-ranked contexts. Current attacks are characterized along two axes: (1) Knowledge accessibility: White-box (full model access) vs. black-box (query-only) approaches. (2) Perturbation granularity: Word-level [raval2020one, liu2023topic, wu2023prada], Phrase-level [song2020adversarial, liu2022order, chen2025flipedrag], Sentence-level [chen2023towards], and Hybrid [liu2024multi].These neural semantic attacks parallel black-hat SEO tactics [gyongyi2005web] but introduce novel challenges. Successful attacks propagate adversarial content to generators, enabling misinformation injection while evading traditional safeguards [zhong2023poisoning, chen2025flipedrag], necessitating integrated security frameworks for RAG systems.
2.3 Attacks to LLMs
@@
Existing attacks on LLMs include jailbreak attacks [wei2024jailbroken, deng2024masterkey, li2023multi, lin2024figure], backdoor attacks [kandpal2023backdoor, lu2024test, cao2023stealthy], prompt injection [liu2023prompt, perez2022ignore, greshake2023not], and poisoning attacks [shafran2024machine, zou2024poisonedrag, cho2024typos, zhong2023poisoning, chen2025flipedrag, zhang2024hijackrag]. Poisoning attacks uniquely threaten RAG systems by injecting adversarial content into retrieval corpora to manipulate outputs. Current RAG poisoning studies focus on closed-domain factoid QA pairs (e.g., "CEO of OpenAI") [zou2024poisonedrag, zhong2023poisoning, cho2024typos, shafran2024machine, zhang2024hijackrag], overlooking the multi-query nature of real-world topic exploration[liu2023topic, xue2024badrag], e.g., "smartwatch battery life" and "health tracking accuracy" under the "wearable tech" theme. This motivates our focus on practical topic-level universal perturbations mirroring universal adversarial examples.
@@
Furthermore, existing defenses based on fact-checking are insufficient for opinion-based queries, such as "Should genetic testing be regulated?", which require nuanced reasoning rather than simple factual recall. Building on the insights from FlippedRAG [chen2025flipedrag], we tackle this significant gap by examining the adversarial manipulation of controversial topics. In such cases, LLMs are required to reason and synthesize multiple perspectives, making them particularly vulnerable to systematic knowledge poisoning.
3 Threat Model
@@
Given a set of topic-queries $Q=\{q_{1},q_{2},\ldots,q_{|Q|}\}$, a RAG corpus $D$, and a desired target polarity $S_{t}$ (e.g., Pro or Con), The adversary’s objective is to steer the RAG system’s overall stance across the entire set of queries toward stance $S_{t}$. Let $\text{doc}_{\text{tar}}$ be the target document the attacker aims to promote. The black-box retrieval model RM assigns a relevance score $R(q,d)$ to each document $d\in D$ and returns the top-$k$ documents as $\text{RM}_{k}(q)=\{d_{1},d_{2},\ldots,d_{k}\}$. The LLM then generates an answer $\text{LLM}(q,RM_{k}(q))$, and the extracted opinion from this response is $S_{o}=\text{S}(\text{LLM}(q,RM_{k}(q)))$.
@@
Figure 2: An overview of our proposed Topic-Orientated Adversarial Opinion Manipulation Attack method for RAG systems.
3.1 Objective of the Adversary
@@
The adversary seeks to modify the target document $doc_{tar}$ through a series of subtle perturbations $p_{adv}$, while preserving its original semantics and fluency, thereby transforming the corpus $D$ into $D^{\prime}=D(\text{doc};\text{doc}_{tar}\oplus p_{\text{adv}})$. By increasing the retrieval model’s relevance score $R(q,\text{doc}_{tar}\oplus p_{\text{adv}})$ for each query $q\in Q$, the promoted document is more likely to appear among the top-$k$ results $\text{RM}_{k}(q)$.
@@
As a result, when the LLM generates answers conditioned on the query and the top-$k$ retrieval results $(q,\text{RM}_{k}(q))$, the adversarially augmented document will influence the opinion $S(\text{LLM}(q,\text{RM}_{k}(q)))$ towards the target polarity $S_{t}$. Crucially, the objective extends beyond a single query: the adversary aims to ensure that, on average, the opinions produced by the LLM across the entire query set $Q$ align with $S_{t}$. In other words, the attacker seeks:
$\max_{p_{\text{adv}}}\frac{1}{|Q|}\sum_{q\in Q}I\bigl(S(\text{LLM}(q,\text{RM}_{k}(q)))=S_{t}\bigr)\hskip 14.22636pt$
@@
where $I(\cdot)$ is an indicator function that returns $1$ if the generated stance matches $S_{t}$ and 0 otherwise. By optimizing for this criterion, the adversary ensures that the aggregated stance across all topic queries shifts significantly toward the desired polarity.
3.2 Capabilities of the Adversary